OpenStudy (daenio):

a) What is the probability that two Data Management and two Calculus students are chosen? b) What is the probability that no Calculus students are chosen? c) What is the probability that all four students are from the Functions class? d) What is the expected number of Data Management student to be selected

OpenStudy (daenio):

a) What is the probability that two Data Management and two Calculus students are chosen? [2] P(x) = (7 C 2) * (4 C 0) * (6 C 2)/ 17 C 4) = 21 * 1 * 15/ 2380 = 315/ 2380 = 0.1324 b) What is the probability that no Calculus students are chosen? [2] P( x=0) = (6 C 0) (13 C 4) / 17 C 4 = 1 * 715/ 2380 = 0.3004 c) What is the probability that all four students are from the Functions class? [2] P (x=4) = (4 C 4) (13 C 0) / 17 C 4 = 1/ 2380 = 0.000420168 d) What is the expected number of Data Management student to be selected? [2] E(x)= 4 x 7/ 17 = 1.65

OpenStudy (daenio):

@jim_thompson5910

OpenStudy (daenio):

Did I do this right?

OpenStudy (daenio):

@imranmeah91 @Ishaan94 @LagrangeSon678

OpenStudy (daenio):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Is there more to the problem that isn't here?

OpenStudy (daenio):

WHOOPS! LOL. Woodsborough High wants to send four Grade 12 students to a Mathematics conference at McGill University. Students with an average of 90% or above are allowed to attend the conference. The Data Management class was able to nominate 7 candidates, the Functions class nominated 4 candidates, and the Calculus class nominated 6 candidates (no student is taking more than one math class). The candidates are chosen at random.

OpenStudy (daenio):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

one moment while I check something

jimthompson5910 (jim_thompson5910):

your answer to part a) is correct

jimthompson5910 (jim_thompson5910):

Everything else checks out, you have all correct answers. Nice work.

OpenStudy (daenio):

Alright thanks! :)

jimthompson5910 (jim_thompson5910):

yw